Problem 5-20 CVP Applications: Break-Even Analysis; Cost Structure; Target Sales [LO5-1, LO5-3, LO5- 4, LO5-5, LO5-6, LO5-8]5. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plReq 4 Req 1 Req 2 Req 3 Req 5 Req 6A Req 6B Due to an increase in labor rates, the company estimates that next years variablReq 6B Req 1 Req 2 Req 3 Req 4 Req 5 Req 6A Refer to the data in Required (2). If the expected change in variable expenses taReq 1 Req 2 Req 3 Req 4 Req 5 Req 6A Req 6B Refer again to the data in Required (2). The president feels that the company musReq 6A Req 6B Req 1 Req 2 Req 3 Req 4 Req 5 Refer to the original data. The company is discussing the construction of a new,If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $225,000, asReq 1 Req 2 Req 3 Req 4 Req 5 Req 6A Req 6B Assume the new plant is built and that next year the company manufactures and sel
Answer #1

Solution 1:
Contribution margin ratio = Contribution margin / sales = $600,000 / $1,500,000 = 40%
Contribution margin per unit = $25 - $15 = $10 per unit
Breakeven sales units = Fixed cost / contribution margin per unit = $375,000 / 10 = 37500 units
Degree of operating leverage = Contribution margin / Net operating income = $600,000 / $225,000 = 2.67


Solution 2:
New variable cost per unit =  $15 + $3 = $18 per ball
new contribution margin per unit = $25 - $18 = $7 per unit
New contribution margin ratio = $7 / $25  =28%
New breakeven point in balls = $375,000 / $7 = 53571 units


Solution 3:
Nos of balls to be sold to earn target income = (Fixed cost + Target profit) / contribution margin per unit
= ($375,000 + $225,000) / $7 = 85714 units

Solution 4:
Variable cost per unit = $18 per unit
Required contribution margin ratio = 40%
required variable cost ratio = 60%
New selling price per unit = $18 / 60% = $30 per unit


Solution 5:
New variable cost per unit = $15 * 60% = $9 per unit
New contribution margin per unit = 25- $9 = $16 per unit
New fixed costs = $375,000*2 = $750,000
New CM ratio = $16/$25 = 64%
New breakeven point = $750,000/ $16 = 46875 units

Solution 6a:
Nos of balls to be sold to earn target income = (Fixed cost + Target profit) / contribution margin per unit
= ($750,000 + $225,000) / $16 = 60938 units

Solution 6b:

E AR M + 6 000 10 11 Northwood company Contribution margin income statement Particulars Amount Sales (60000*$25) 1500000 Vari